OpenStudy (pottersheep):

Grade 12 vectors help please! Intersection of planes For what value of k will the following set of planes intersect in a line? x-2y-z=0 x+9y-5z=0 kx-y+z=0 I have to solve this using matrices. Thank you

OpenStudy (anonymous):

x-2y-z=0 x+9y-5z=0 kx-y+z=0 it has a solution if 1{(9)(1)-(-5)(-1)}+2{(1)(1)-(-5)(k)}-1{(1)(-1)-(9)(k)}=0 1(9-5)+2(1+5k)-1(-1-9k)=0 4+2+10k+1+9k=0 19k=7 |■(1&-2&-1@1&9&-5@k&-1&1)|=0 K=7/19

OpenStudy (pottersheep):

Thank you!!!!